# Extracting and Transferring Abilities For Building Multi-lingual Ability-enhanced Large Language Models

### v1

@@ 4.3 The Overall Procedure
To better demonstrate our approach, we present the complete procedure in Algorithm 1 in the pseudo-code form. The procedure of MAET consists of two main stages, i.e., ability-related weights extraction and multi-lingual ability transferring. For the extraction stage, we first utilize the accumulated gradient to estimate the importance of each neuron by Equation 1. Then, we leverage the model trained on the general corpus to remove the influence of language and obtain the ability-specific weight through Equation 2. In the transfer stage, we utilize Equation 3 and Equation 4 to obtain the multi-lingual weight and identify the ability-related parameter tensors in LLM. After it, we leverage Equation 5 to fulfill the multi-lingual ability transfer, to build the multi-lingual ability-enhanced LLM.
@@ 5.1 Experimental Settings
In this part, we introduce the details of our experimental settings, including the datasets utilized in the continual pre-training and evaluation process, baseline methods for comparison, and the implementation details of our approach.
@@
Datasets. In this work, we focus on transferring the advanced abilities (i.e., mathematical and scientific reasoning abilities) of LLMs from English scenarios to multi-lingual scenarios, including high-resource languages (i.e., Spanish and Chinese) and low-resource languages (i.e., Bengali and Telugu). For the training corpus, we extract the corpus of the corresponding languages from the dataset proposed by previous work (Yang et al., 2023; Scao et al., 2022; Laurençon et al., 2022) as the general training corpus, and utilize OpenWebMath (Paster et al., 2024) and the arXiv papers (Soldaini et al., 2024) as the ability-related corpus for mathematical tasks and scientific tasks respectively. For the evaluation benchmark, we follow the evaluation settings in previous work (OpenAI, 2023), utilizing Multi-lingual Grade School Math (MGSM) (Shi et al., 2023) and science tasks from multi-lingual MMLU (Lai et al., 2023) (i.e., college biology, college chemistry, college physics, high school biology, high school chemistry, and high school physics) as the downstream tasks for multi-lingual scenarios. The statistical information of the datasets is presented in Table 6.
@@
Baselines. In our evaluation, a baseline can be represented as three parts, i.e., training parameters, training approach, and training data. First, we conduct the full parameters training and the LoRA training (Hu et al., 2022) in our evaluation, denoted as the “F” and “L” at the prefix of the training approaches, respectively. For the training approach, we employ continual pre-training (CPT) (Gururangan et al., 2020), domain adaption (DA) (Taylor et al., 2022), and model merging with task vector (TV) (Ilharco et al., 2023). Besides, for the training data, “L” and “A” refer to the multi-lingual general corpus and the single-lingual ability-related corpus, respectively. Moreover, to conduct a more comprehensive evaluation, we also present the performance of different LLMs, i.e., Baichuan-2 7B (Yang et al., 2023), Mistral 7B (Jiang et al., 2023), LLaMA-2 7B (Touvron et al., 2023), and LLaMA-3 8B (Dubey et al., 2024).
Implementation Details. In the experiment, we adapt LLaMA-3 8B as the backbone LLM, and employ Transformers (Wolf et al., 2020) and Deepspeed  framework to perform the training process. For the training process, the learning rate, batch size, and training step are set as $5\times 10^{-5}$, $1$M tokens, and $2$B tokens, respectively. Besides, for the key neurons locating, we select the top 5% relevant neurons as the key neuron set $\mathcal{N}$ for both stages and identify the last 80% and 60% similar tensor as the key sub-network $\mathcal{T}$ for mathematical reasoning tasks and scientific reasoning tasks respectively. We present model details about hyper-parameters in Appendix B.
5.2 Main Results
To comprehensively evaluate our proposed MAET, we employ MAET on mathematical and scientific tasks in multi-lingual scenarios and present the results in Table 2.
[…]
Methods Multilingual Mathematical Tasks Multilingual Scientific Tasks
ES ZH BN TE Avg. ES ZH BN TE Avg.
Baichuan-2 7B 17.20 28.00 4.80 2.40 13.10 42.27 46.43 30.17 26.21 36.27
Mistral 7B 38.80 34.40 9.60 2.80 21.40 52.08 45.33 32.91 27.96 39.57
LLaMA-2 7B 7.60 12.00 1.60 0.00 5.30 34.16 31.68 24.56 22.15 28.14
LLaMA-3 8B 48.40 38.80 28.80 20.40 34.10 55.06 47.24 36.63 29.26 42.05
@@ 5.2 Main Results
+ F-TV 42.00 32.40 16.00 10.40 25.20 53.36 46.57 36.70 30.73 41.84
+ L-TV 45.60 39.20 30.80 25.60 35.30 55.46 48.27 36.65 30.44 42.71
+ MAET (Ours) 49.60 41.60 32.40 25.20 37.20 56.20 48.00 37.64 30.38 43.06
Table 2: The performance comparison of different approaches on multilingual mathematical tasks and multilingual scientific tasks. Avg. denotes the average accuracy of the multi-lingual tasks. ES, ZH, BN, and TE denote Spanish, Chinese, Bengali, and Telugu, respectively. The best is in bold and the second best is underlined.
Figure 2: The results of ablation study. KNL, AW, API, and AT denote key neurons locating (Eq. 1), ability weights obtaining (Eq. 2), ability-related parameter tensors identifying (Eq. 4), and advanced abilities transferring (Eq. 5).
@@
5.3 Detailed Analysis
To comprehensively evaluate our proposed approach MAET and analyze its features, we conduct several experiments and detailed analysis in this part, including the ablation study, analysis of the transfer ratio of LLM parameters, and the generalization of MAET.
Ratio of Key Parameters During Transferring Stage. Identifying and updating the ability-related sub-network of LLMs is the key point of our MAET. We conduct experiments to analyze the influence of the transferring ratio $k_{2}\%$ and show the results in Figure 3. Observing the results, the performance of LLM has decreased in a lower and higher ratio of the ability-related parameters identifying process. The main reason is that the lower proportion transfers incomplete knowledge to the model and makes LLM unable to possess the corresponding ability, affecting the performance on the downstream tasks. In contrast, the higher proportion cannot extract the ability weight precisely and will transfer too much language-related knowledge to the model, making the conflict with the LLM inner knowledge and hurting the multi-lingual scenarios performance.
@@ 5.3 Detailed Analysis
Methods MMLU HumanEval MBPP OpenbookQA
LLaMA-3 8B 60.85 35.98 46.60 65.00
+ CPT 58.46 (-2.39) 28.66 (-7.32) 39.60 (-7.00) 61.40 (-3.60)
+ MAET (Ours) 61.07 (+0.22) 35.98 (+0.00) 47.40 (+0.80) 65.00 (+0.00)
Table 4: The out-of-domain performance comparison of different training methods to train LLaMA-3 8B on OpenWebMath. During the ability-enhancing process, previous methods will hurt the OOD abilities of LLM, while our MAET can maintain the corresponding abilities.
@@
Ability-related Sub-networks of LLM. To assess and probe the ability-related sub-networks of LLMs, we only transfer the specific tensors (i.e., tensors in self-attention and MLP mechanism) from the ability weight to the final models through Eq. 5, to analyze the LLM inner abilities. The experimental results are presented in Table 3. From the experiment, we can observe that although the proportion of MLP layers (41.38%) is lower than the attention layers (45.26%), only transferring the MLP layers outperforms transferring the attention layers, indicating that the MLP layers are more related to the advanced abilities and stores the corresponding knowledge. In the MLP layers of LLM, the gate mechanism (i.e., MLP Gate) will control the transmission of information and the down project mechanism (i.e., MLP Down) will integrate the knowledge from previous layers, so that transferring the MLP layers can achieve better performance on the downstream tasks.
Out-of-Domain performance of MAET. We conduct experiments about adapting mathematical ability on the general LLM through MAET, and assess the performance on out-of-domain (OOD) tasks (i.e., MMLU (Hendrycks et al., 2021), HumanEval (Chen et al., 2021a), MBPP (Austin et al., 2021), and OpenbookQA (Mihaylov et al., 2018)), which can reflect and assess different abilities of LLMs. Results are presented in Table 4. We can observe that the performance of LLM on all evaluation tasks has decreased through the CPT training process, and the maximum decrease has been achieved 7.32% on the HumanEval task. One of the possible reasons is that LLaMA-3 has been trained on OpenWebMath during pre-training and the CPT process makes it overfit and forget the knowledge of other domains, hurting the performance on OOD tasks. In contrast, our proposed MAET achieves comparable and even better performance with backbone LLM in all downstream scenarios. Since we identify and update the key neurons related to the specific ability, the ability of LLM can be precisely enhanced, and this strategy also helps the OOD tasks needed for mathematical ability, e.g., mathematical sub-tasks in MMLU and code synthesis task MBPP.
6 Conclusion
7 Limitations
In this section, we discuss the limitations of our work. First, we only implement our approach MAET on 8B LLMs (i.e., LLaMA-3 8B), and do not adopt the LLMs with larger scales (e.g., 13B or 70B LLMs) in the experiment, due to the limitation of computational resources. We will test the effectiveness of our approach on these LLMs in the future. Second, we only evaluate our approach on two downstream tasks (i.e., mathematical and scientific reasoning tasks) in multi-lingual scenarios. Although they are challenging and widely-used testbeds, it is still meaningful to verify our methods on other tasks. Whereas, as we test the performance on diverse high-resource and low-resource languages, it can also provide comprehensive performance estimation for our approach in multi-lingual scenarios. Third, our approach is a general method for ability transferring across different domains, but in this work, we only consider the multi-lingual scenarios and obtain a multi-lingual LLM with the specific ability being enhanced. Forth, our approach also relies on continual pre-training the LLM. Although the training corpus is not very large, it also increases the cost. Fortunately, after we obtain the pre-trained weights, our following steps only need simple addition and subtraction operations for ability transferring, which is flexible for online deployment and application. In future work, we will focus on reducing the data requirement for the pre-training corpus and also improving the weights extracting efficiency. Finally, we do not consider the potential risk and ethics issues that might hurt the alignment of LLMs when using our approach. Actually, our approach is also applicable to transfer the alignment ability across languages. We will investigate to it in the future.
Appendix A Empirical Study
Figure 4: The results of empirical experiments. We present the loss of different training methods during the training process, the cosine similarity of LLM layers after being trained on Zhihu and Reddit, and the similarity of LLMs being trained on different training corpus.
Appendix B Details of Hyper-parameters
We release all of the hyper-parameters during our experiment to better reproduce our proposed approach. Table 5 shows the details of hyper-parameters of different stages.
Stage Hyper-Parameter Mathematical Tasks Scientific Tasks
Extraction Learning Rate $5\times 10^{-5}$ $5\times 10^{-5}$
Batch Size 1M Tokens 1M Tokens
Training Steps 2B Tokens 2B Tokens
$\alpha$ in Extraction 0.8 0.8
$\beta$ in Extraction 0.2 0.2
Ratio of Key Neurons 5% 5%
Extraction Learning Rate $5\times 10^{-5}$ $5\times 10^{-5}$
Batch Size 1M Tokens 1M Tokens
Training Steps 2B Tokens 2B Tokens
$\gamma$ in Transferring 0.2 0.2
$\eta$ in Transferring 1.0 1.0
Ratio of Key Neurons 80% 60%
Table 5: The details of hyper-parameters in the evaluation.
Appendix C Details of Dataset
Language Training Dataset (Tokens) Evaluation Dataset (Instances)
English 1.81B 1.30B (Math) / 1.82B (Sci) 250 1,245
Spanish 1.81B - 250 1,232
Chinese 1.80B - 250 1,229
Bengali 1.81B - 250 1,137
Telugu 1.81B - 250 1,036

### v2

@@ 4.3 The Overall Procedure
To better demonstrate our approach, we present key concepts in Table 4 for further clarifying and provide the complete procedure in Algorithm 1 in the pseudo-code form. The procedure of MAET consists of two main stages, i.e., ability-related weights extraction and multi-lingual ability transferring. For the extraction stage, we first utilize the accumulated gradient to estimate the importance of each neuron by Equation 1. Then, we leverage the model trained on the general corpus to remove the influence of language and obtain the ability-related weight through Equation 2. In the transfer stage, we utilize Equation 3 and Equation 4 to obtain the multi-lingual weight and identify the ability-related parameter tensors in LLM. After it, we leverage Equation 5 to fulfill the multi-lingual ability transfer, to build the multi-lingual ability-enhanced LLM.
@@ 5.1 Experimental Settings
We introduce the datasets and the baseline methods in our evaluation in this part, and present the implementation details of our approach in Appendix B.
@@
Datasets. We focus on transferring the advanced abilities (i.e., mathematical and scientific reasoning abilities) of LLMs from English scenarios to multi-lingual scenarios, including high-resource languages (i.e., Spanish) and low-resource languages (i.e., Bengali and Telugu). Thus, for the training corpus, we extract the corpus proposed by previous work (Yang et al., 2023; Scao et al., 2022; Laurençon et al., 2022) as the general training corpus, and utilize OpenWebMath (Paster et al., 2024) and the arXiv papers (Soldaini et al., 2024) as the ability-related corpus for mathematical tasks and scientific tasks, respectively. For the evaluation benchmark, we follow the evaluation settings in previous work (OpenAI, 2023), utilizing Multi-lingual Grade School Math (MGSM) (Shi et al., 2023) and science tasks from multi-lingual MMLU (Lai et al., 2023) (i.e., college biology, college chemistry, college physics, high school biology, high school chemistry, and high school physics) as the downstream tasks. The statistical information of the datasets is presented in Table 6 in the Appendix.
@@
Baselines. In our evaluation, we utilize LLaMA-3 8B (Dubey et al., 2024) as the backbone model and adopt four categories of the widely used methods as baselines, i.e., continually pre-training, transfer learning, data augmentation, and model merging based approaches. Concretely, a baseline can be represented as three parts, i.e., training parameters, training approach, and training data. First, we conduct the full parameters training and the LoRA training (Hu et al., 2022), which are denoted as the “F” and “L” at the prefix, respectively. Second, for the training approach, we employ continual pre-training (CPT) (Gururangan et al., 2020), domain adaption (DA) (Taylor et al., 2022), and model merging with task vector (TV) (Ilharco et al., 2023). Third, for the training data, “L”, “A”, and “T” refer to the multi-lingual general corpus, English ability-related corpus, and multi-lingual ability-related corpus translated from GPT-4o (Hurst et al., 2024), respectively. Moreover, we also present the performance of other LLMs, i.e., Baichuan-2 7B (Yang et al., 2023), Mistral 7B (Jiang et al., 2023), and LLaMA-2 7B (Touvron et al., 2023).
Methods #Tokens Multilingual Mathematical Tasks Multilingual Scientific Tasks
ES BN TE Avg. RICER ($\uparrow$) ES BN TE Avg. RICER ($\uparrow$)
Baichuan-2 7B - 17.2 4.8 2.4 8.1 - 42.3 30.2 26.2 32.9 -
Mistral 7B - 38.8 9.6 2.8 17.1 - 52.1 32.9 28.0 37.7 -
LLaMA-2 7B - 7.6 1.6 0.0 3.1 - 34.2 24.6 22.2 27.0 -
LLaMA-3 8B - 48.4 28.8 20.4 32.5 - 55.1 36.6 29.3 40.3 -
Vanilla Continually Pre-training based Approaches
@@
[…]
Data Augmentation based Approaches
[…]
Model Merging based Approaches
+ F-TV 12B 42.0 16.0 10.4 22.8 -80.8 53.4 36.7 30.7 40.3 -0.3
+ L-TV 12B 45.6 30.8 25.6 34.0 12.5 55.5 36.7 30.4 40.9 4.6
+ MAET (Ours) 12B 49.6 32.4 25.2 35.7 26.9 56.2 37.6 30.4 41.4 9.2
Table 2: The performance of different approaches on multilingual mathematical and scientific tasks. ES, BN, and TE denote Spanish, Bengali, and Telugu, respectively. #Tokens denotes the number of training tokens. RICER refers to the reciprocal of incremental cost effectiveness ratio, i.e., RICER = #Tokens / Improvement ×$\times 100\%$.
5.2 Main Results
We present the evaluation results of MAET and the baseline methods in Table 2.
[…]
Second, although our MAET shows similar training efficiency to transfer learning based approaches, MAET performs better than these baseline methods, showing the higher RICER score (e.g., 26.9 v.s. 16.9). For transfer learning based approaches, since the model is only trained on the single-lingual ability-related corpus during the domain adaptation process, it is difficult for LLM to handle the challenging tasks in multi-lingual scenarios. Concretely, the performance of LLaMA-3 8B on the multi-lingual scientific tasks even decreases after domain adaptation, showing a 4% relative decrease. To alleviate this issue, MAET leverages the calculation between the parameters of different models to extract the ability-related weights, which are language-agnostic and can be transferred to any other scenario.
Third, MAET also achieves higher performance than data augmentation based approaches (i.e., training LLM on the multi-lingual ability-related corpus translated by GPT-4o). The translation-based method consumes more computational resources and cannot achieve better performance. The reason might be that LLMs cannot perform the translation process well and the translated corpus shares similar knowledge of the specific domain, which makes LLM overfit the corresponding knowledge and cannot really understand the specific knowledge. In contrast, our approach decomposes the advanced ability and language ability, and transfers the advanced ability from one language to another, preventing overfitting, decreasing the expense, and improving performance. These results demonstrate that data-centric methods are difficult to build a multi-lingual ability-enhanced LLM.
5.3 Detailed Analysis
To further analyze the features of MAET, we conduct the detailed analysis in this part, including the ablation study, analysis of the transfer ratio $k_{2}$, and the generalization of MAET.
Figure 2: The ablation study. KNL, AWO, ATI, and AAT denote key neurons locating (Eq. 1), ability weights obtaining (Eq. 2), ability-related tensors identifying (Eq. 4), and advanced abilities transferring (Eq. 5).
Ratio of Key Parameters During Transferring Stage. Identifying and updating the ability-related sub-network of LLMs is the key point of our MAET. We analyze the influence of the transferring ratio $k_{2}\%$ and show the results in Figure 3. Observing the results, the performance of LLM has decreased in a lower and higher ratio of the ability-related parameters identifying process. The main reason is that the lower proportion transfers incomplete knowledge to the model and makes LLM unable to possess the corresponding ability, affecting the performance on the downstream tasks. In contrast, the higher proportion cannot extract the ability weight precisely and will transfer too much language-related knowledge to the model, making the conflict with the LLM’s inner knowledge and hurting the multi-lingual scenarios’ performance.
@@ 5.3 Detailed Analysis
Out-of-Domain Performance of MAET. We conduct experiments about adapting mathematical ability on LLaMA-3 8B through MAET, and assess its performance on out-of-domain (OOD) tasks (i.e., MMLU (Hendrycks et al., 2021), HumanEval (Chen et al., 2021a), MBPP (Austin et al., 2021), and OpenbookQA (Mihaylov et al., 2018)). Results are presented in Table 3. We can observe that the performance of LLM on all evaluation tasks has decreased through the CPT training process, and the maximum decrease has been achieved 7.32% on the HumanEval task. One of the possible reasons is that LLaMA-3 has been trained on OpenWebMath during pre-training and the CPT process makes it overfit and forget the knowledge of other domains, hurting the performance on OOD tasks. In contrast, our proposed MAET achieves comparable and even better performance with backbone LLM in all downstream scenarios. Since we identify and update the key neurons related to the specific ability, the ability of LLM can be precisely enhanced, and this strategy also helps the OOD tasks needed for mathematical ability, e.g., mathematical tasks in MMLU and MBPP.
Methods MMLU MBPP OpenbookQA
LLaMA-3 8B 60.85 46.60 65.00
+ CPT -2.39 -7.00 -3.60
+ MAET +0.22 +0.80 +0.00
Table 3: The out-of-domain performance of different training methods to train LLaMA-3 8B on OpenWebMath. During the ability-enhancing process, previous methods will hurt the OOD abilities of LLM, while our MAET can maintain the corresponding abilities.
6 Conclusion
Limitations
In this section, we discuss the limitations of our work. First, we only implement our approach MAET on 8B LLMs (i.e., LLaMA-3 8B), and do not adopt the LLMs with larger scales (e.g., 13B or 70B LLMs) in the experiment, due to the limitation of computational resources. We will test the effectiveness of our approach on these LLMs in the future. Second, we only evaluate our approach on two downstream tasks (i.e., mathematical and scientific reasoning tasks) in multi-lingual scenarios. Although they are challenging and widely-used testbeds, it is still meaningful to verify our methods on other tasks. Whereas, as we test the performance on diverse high-resource and low-resource languages, it can also provide comprehensive performance estimation for our approach in multi-lingual scenarios. Finally, we do not consider the potential risk and ethics issues that might hurt the alignment of LLMs when using our approach. Actually, our approach is also applicable to transfer the alignment ability across languages. We will investigate to it in the future.
Appendix A Empirical Study
Stage Hyper-Parameter Mathematical Tasks Scientific Tasks
Extracting Learning Rate $5\times 10^{-5}$ $5\times 10^{-5}$
Batch Size 1M Tokens 1M Tokens
Training Steps 2B Tokens 2B Tokens
$\alpha$ in Extraction 0.8 0.8
$\beta$ in Extraction 0.2 0.2
Ratio of Key Neurons 5% 5%
Transferring Learning Rate $5\times 10^{-5}$ $5\times 10^{-5}$
Batch Size 1M Tokens 1M Tokens
Training Steps 2B Tokens 2B Tokens
$\gamma$ in Transferring 0.2 0.2
$\eta$ in Transferring 1.0 1.0
Ratio of Key Neurons 80% 60%
$\mu$ for Spanish 1.5 1.5
$\mu$ for Chinese 2.0 2.0
$\mu$ for Bengali 1.2 1.2
$\mu$ for Telugu 1.2 1.2
Table 5: The details of hyper-parameters in the training and evaluation process.
Language Training Dataset (Tokens) Evaluation Dataset (Instances)
English 1.81B 1.30B (Math) / 1.82B (Sci) 250 1,245
Spanish 1.81B - 250 1,232
Chinese 1.80B - 250 1,229
Bengali 1.81B - 250 1,137
Telugu 1.81B - 250 1,036
Methods Qwen2.5 0.5B Gemma2 2B
ES TE Avg. ES TE Avg.
Backbone LLM 36.64 25.69 31.17 43.41 30.01 36.71
[…]
+ MAET w/o API 36.72 28.91 32.82 43.23 29.59 36.41
+ MAET (Ours) 36.91 29.62 33.27 43.62 30.37 37.00
Table 7: The performance comparison of different LLMs on multilingual scientific tasks.
Appendix B Implementation Details
In the experiment, we adapt LLaMA-3 8B as the backbone LLM, and employ Transformers (Wolf et al., 2020) and Deepspeed  framework to perform the training process. And we also present the evaluation results of different backbone LLM (i.e., Qwen2.5 0.5B (Hui et al., 2024) and Gemma2 2B (Rivière et al., 2024)) in Appendix E. For the training process, the learning rate, batch size, and training step are set as $5\times 10^{-5}$, $1$M tokens, and $2$B tokens, respectively. Besides, for the key neurons locating, we select the top 5% relevant neurons as the key neuron set $\mathcal{N}$ for both stages and identify the last 80% and 60% similar tensor as the key sub-network $\mathcal{T}$ for mathematical reasoning tasks and scientific reasoning tasks respectively.
Hyper-parameters Selection. we released all of the hyper-parameters during our experiment in Table 5, to reproduce our proposed approach better. The hyperparameters discussed in the paper can be categorized into two types: training-related parameters (e.g., learning rate, batch size) and training-independent parameters (i.e., $\alpha$, $\beta$, $\gamma$, $\eta$, and $\mu$). Training-related parameters do not require extensive hyperparameter tuning, as existing studies Dubey et al. (2024); Hui et al. (2024) provide clear guidelines for setting them. On the other hand, training-independent parameters are used to construct ability-related weights, tensors, and language-specific weights. These techniques are similar to those employed in model merging Ilharco et al. (2023); Yadav et al. (2023), and the hyperparameter setting approach outlined in the paper can be applied. A limited number of hyperparameter sets can be defined and validated, as the process primarily involves simple additions and subtractions of model parameters, making it computationally inexpensive.
Appendix C Details of Dataset
Appendix D Prompt for Translation
You should translate the following text from English to {TARGET LANGUAGE} and should not modify the latex code or website code. You should not add any details that are not mentioned in the original text.
## English
{ENGLISH TEXT}
## {TARGET LANGUAGE}
Appendix E Performance of Small Scale LLMs
We conduct the different LLMs with different sizes (i.e.,, Qwen2.5-0.5B and Gemma2-2B) in our experiment to valid the practicality of our approach. We assess MAET and baselines on multi-lingual scientific reasoning tasks and present the evaluation results in Table 7. Comparing the performance of MAET and the baseline methods, we can observe that MAET can also enhance the performance of small scale models and outperform competitive baselines. Therefore, the evaluation results have shown the effectiveness of MAET and verified that MAET is a general LLM enhancement technology.
@@ Appendix E Performance of Small Scale LLMs
Appendix F Ability-related Sub-networks of LLM
@@
To assess and probe the ability-related sub-networks of LLMs, we only transfer the specific tensors (i.e., tensors in self-attention and MLP mechanism) from the ability weight to the final models through Eq. 5, to analyze the LLM inner abilities. The experimental results are presented in Table 8. From the experiment, we can observe that although the proportion of MLP layers (41.38%) is lower than the attention layers (45.26%), only transferring the MLP layers outperforms transferring the attention layers, indicating that the MLP layers are more related to the advanced abilities and stores the corresponding knowledge. In the MLP layers of LLM, the gate mechanism (i.e., MLP Gate) will control the transmission of information and the down project mechanism (i.e., MLP Down) will integrate the knowledge from previous layers, so that transferring the MLP layers can achieve better performance on the downstream tasks.